Circular No. 6134

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995D IN NGC 2962 S. Nakano, Sumoto, Japan, reports the discovery by Reiki Kushida

(Yatsugatake South Base Observatory) of an apparent supernova of mag 14.0 on an unfiltered CCD frame taken on Feb. 10.756 UT with a 0.25-m f/6.3 Schmidt-Cassegrain telescope. Y. Kushida has measured the following accurate position from five CCD images by the discoverer: α = 9^{h}40^{m}54^{s}.79, δ = +5^{o}08′26′′.6 (equinox 2000.0), which corresponds to a point 11′′ east and 90′′.5 south of the center of NGC 2962; no image brighter than mag 17 appears at this position on CCD images secured by R. Kushida on 1994 Dec. 3, on 1995 Jan. 13, and on Jan. 25. Comet patrol T-Max films taken by Y. Kushida on Jan. 31 show no star brighter than mag 14.8. A further unfiltered CCD image by Y. Kushida on Feb. 11.475 shows SN 1995D at mag 13.9 ± 0.2.

NOVA AQUILAE 1995 R. M. Wagner and R. Bertram, Ohio State University; and S. G. Starr-

field, Arizona State University, report: “Optical spectra (range 450–570 nm, resolution 1 nm) of the nova candidate described on IAUC 6133 were obtained at the Perkins 1.8-m telescope on Feb. 10.54 UT. The spectra show strong Hβ and permitted Fe II emission lines. The continuum over this spectral range is relatively flat. The FWHM of Hβ emission is ∼ 1200 km/s. The spectra confirm this object as a slow ‘Fe II’-class nova in its post-maximum phase of development.”

O. Ohshima, M. Shimizu, and K. Ayani, Bisei Astronomical Observatory; and M. Suzuki, Kanazawa Institute of Technology, communicate: “We obtained spectra of this object with the Bisei 1.01-m reflector (+ CCD spectrograph) on Feb. 10.9 UT. Strong emission lines of Hα, Hβ, Fe II (492.3, 501.8, 516.9 nm), [N II] (575.4 nm), and Na I (589.0 nm) were identified clearly. The FWHM of the Hα line (corrected by the instrumental FWHM of 1.18 nm) is 2.94 nm, which corresponds to a velocity dispersion of 1340 km/s. The Hα profile shows red-wing asymmetry.”

Additional visual magnitude estimates: Feb. 10.18 UT, 8.8 (L. Kiss, Szeged, Hungary); 10.198, 8.9 (L. Szentaskó, Veresegyház, Hungary); 10.551, 8.6 (P. Collins, Scottsdale, AZ).

1995 February 11 Daniel W. E. Green